# Exact Sequence Interpolation with Transformers

### v2

@@ 5.1. Proof Strategy
Our proof proceeds in four steps, illustrated in Figure˜5 for a simple sequence interpolation problem in $\mathbb{R}^{2}$ with $N=3$ input sequences $X^{1}$, $X^{2}$ and $X^{3}$ of length $n^{1}=5$, $n^{2}=3$ and $n^{3}=3$ and output sequences of length $m^{1}=1,m^{2}=1$, and $m^{3}=2$. The initial configuration of the sequences and the desired output sequences are shown in panel $(a)$ of Figure˜5.
@@
Figure 5. Proof sketch of Theorem˜1.4 applied to $N=3$ input sequences in $\mathbb{R}^{2}$ (denoted with blue circles, orange triangles and green stars, respectively), with output sequences of length $m^{1}=1,m^{2}=1$, and $m^{3}=2$ (denoted with diamonds colored accordingly). The tokens of the initial sequences are shown in panel $(a)$. Panels (b)(b)$(e)$ show the tokens of each sequence after the separation, leader selection, collapse and interpolation steps of the proof. Note that, after the collapse step in panel $(d)$, only tokens selected as leaders are visible. After all steps, sequences match their outputs, as shown in panel $(e)$.
  1. (1)
@@
    Separation. In this step, we use $\mathcal{O}(N)$ transformer blocks to separate overlapping tokens from different input sequences (see panel $(b)$ in Figure˜5). This step requires alternating feed-forward and self-attention layers.
  2. (2)
@@
    Leader selection. In this step, we move all tokens to the positive orthant. Then, we select $m^{j}$ points from each sequence to act as “leaders”, and position them carefully using $\mathcal{O}(\sum_{j=1}^{N}m^{j})$ transformer blocks (see panel $(c)$ in Figure˜5).
  3. (3)
@@
    Collapse. In this step, all tokens not selected to be leaders collapse onto the leaders by leveraging the clustering property of a single transformer block (see panel $(d)$ in Figure˜5).
  4. (4)
@@
    Interpolation. In this final step, the collapsed sequences are driven sequentially to their outputs (see panel $(e)$ in Figure˜5), which requires at most $\mathcal{O}(\sum_{j=1}^{N}m^{j})$ transformer blocks.
@@
The next subsections make these steps rigorous to prove Theorem˜1.4. Throughout, we use the crucial fact that the ReLU is zero in a half-space, allowing us to move some tokens while leaving the other ones fixed.
5.2. Separation
@@
We begin by ensuring that all input tokens are mapped to non-zero vectors, which can be easily obtained with a single transformer block globally shifting all tokens (setting $U=0$, $b=1$ and $W=w\in\mathbb{R}^{d}$ for any non-zero vector in the FF layer, and $\rho=1$, $V=0$, and $A=0$ in the SA layer). After this initial block, we prove the result constructively by induction on $N$ assuming non-zero tokens. Throughout, $x_{i}^{j}$ denotes the $i$-th token of sequence $X^{j}$.
$\bm{N=2}$.
@@
By relabeling tokens if necessary, we may assume without loss of generality that $x_{1}^{2}\in X^{2}\setminus X^{1}$, because the sequences are assumed pairwise distinct. We consider two cases, each of which can be handled using one transformer block.
@@
  • Case 1: $x_{1}^{2}\in\operatorname{ext}\operatorname{\overline{co}}(X^{2})$. In this case, the extreme token $x_{1}^{2}\in X^{2}$ allows for sequence separation using a single transformer block of width $d^{\prime}=1$. For the $\operatorname{\mathrm{FF}}$ layer of this block, we apply Lemma˜4.7 with $i^{\star}=j^{\star}=1$ to construct a feed-forward layer $\mathbb{FF}$ and a vector $v\in\mathbb{R}^{d}$ such that, for some $i_{1}\in[m^{1}]$,
    (5.1a) $\mathcal{C}_{r}(\mathbb{FF}(X^{1}),vv^{\top})=\{i_{1}\}\quad\forall r\in[m^{1}],$
    (5.1b) $\mathcal{C}_{\ell}(\mathbb{FF}(X^{2}),vv^{\top})=\{1\}\quad\forall\ell\in[m^{2}].$
@@
    Note that $x_{i_{1}}^{1}\in\operatorname{ext}\operatorname{\overline{co}}(X^{1})$ by Lemma˜4.5, and $x_{i_{1}}^{1}\neq x_{1}^{2}$ by construction. For the SA layer, instead, we take $0<\alpha<1$ to be specified precisely below and set $\rho=1-\alpha$, $V=\alpha I$ and $A=vv^{\top}$ to obtain
    $\operatorname{\mathrm{SA}}_{r}(\mathbb{FF}(X^{1}))$ $={x}_{r}^{1}+\alpha({x}_{i_{1}}^{1}-{x}_{r}^{1})+b\frac{v}{\|v\|^{2}},$
    $\operatorname{\mathrm{SA}}_{\ell}(\mathbb{FF}(X^{2}))$ $={x}_{\ell}^{2}+\alpha({x}_{1}^{2}-{x}_{\ell}^{2})+b\frac{v}{\|v\|^{2}}.$
@@
    (The last term in these two equations comes from the shift introduced by the FF layer.) These quantities are different for all $\{r,\ell\}\in[m^{1}]\times[m^{2}]$ if and only if
    (5.2) $(1-\alpha)(x_{\ell}^{2}-x_{r}^{1})+\alpha(x_{1}^{2}-x_{i_{1}}^{1})\neq 0.$
@@
    This condition fails if and only if there exists one pair $\{r,\ell\}$ and $c_{r\ell}\in\mathbb{R}$ such that $x_{\ell}^{2}-x_{r}^{1}=c_{r\ell}(x_{1}^{2}-x_{i_{1}}^{1})$ and $\alpha$ is the unique solution to $\frac{\alpha}{\alpha-1}=c_{r\ell}$. This means there are at most finitely many choices of $\alpha$ for which (5.2) fails, so it suffices to pick $\alpha$ not from this set.
@@
    Figure 6. Illustration of Case 2 in Lemma˜5.1 for $N=2$ sequences in $\mathbb{R}^{2}$. The initial tokens in $X^{1}$ (circles) and $X^{2}$ (stars) are shown in panel $(a)$, as well as the $\operatorname{\mathrm{FF}}$ layer that globally shifts by $v\in\mathbb{R}^{2}$ and moves the distinct token from $X^{2}$ (marked with a red arrow). In panel $(b)$, we show the result of applying the resulting FF layer of width $d^{\prime}=4$. In panel $(c)$, the sequences have been separated by the $\operatorname{\mathrm{SA}}$ layer.
@@
    Fix any $y\notin(\operatorname{\overline{co}}(X^{2})\cup X^{1})$. By Lemma˜4.8, there exists a width $d^{\prime}=3$ feed-forward layer $\tilde{\operatorname{\mathrm{FF}}}$ such that $\tilde{\operatorname{\mathrm{FF}}}(x_{1}^{2})=y$, and $\tilde{\operatorname{\mathrm{FF}}}(x)=x$ for all other tokens $x\neq x_{1}^{2}$. Concatenating $\tilde{\operatorname{\mathrm{FF}}}$ to the width $d^{\prime}=1$ feed-forward layer of Lemma˜4.5, we obtain a width $d^{\prime}=4$ feed-forward layer $\mathbb{FF}$ such that ˜5.1 holds and
    (5.3) $\operatorname{\mathrm{FF}}(x_{1}^{2})\in\operatorname{ext}\operatorname{\overline{co}}\mathbb{FF}(X^{2}),$
    since ˜5.3 is invariant under constant shifts. Exactly as in Case 1, the SA layer in this block is given by Lemma˜4.5 (see Figure˜6 for an illustration of both cases).
Induction step.
Assume there exists a hardmax transformer $\operatorname{\mathrm{T}}^{N-1}$ with width $d^{\prime}\leq 4$ and at most 1+2(N2)1+2(N-2) blocks such that
@@
Set $\hat{X}\coloneqq\operatorname{\mathrm{T}}^{N-1}(X)$ to ease the notation. Here and throughout, we write $\operatorname{\mathrm{T}}_{i}(X)$ for the $i$-th token of the transformed sequence $\operatorname{\mathrm{T}}(X)=\{\operatorname{\mathrm{T}}_{1}(X),\dots,\operatorname{\mathrm{T}}_{\operatorname{len}(X)}(X)\}$. Fix without loss of generality $\hat{x}_{1}^{N}\in\hat{X}^{N}$. Our next objective is to build a hardmax transformer $\operatorname{\mathrm{T}}^{\prime}$ with at most two blocks such that
(5.4) $\operatorname{\mathrm{T}}^{\prime}(\hat{X}^{N})\cap\operatorname{\mathrm{T}}^{\prime}(\hat{X}^{j})=\emptyset\quad\text{for all}\;j\in[N-1].$
If $\{\hat{x}_{1}^{N}\}\cap\hat{X}^{j}=\emptyset$ for all $j\in[N-1]$, then we apply exactly the same argument as in Case 1 or Case 2 above for $N=2$, depending on whether $\hat{x}_{1}^{N}\in\operatorname{ext}\operatorname{\overline{co}}(\hat{X}^{N})$ or $\hat{x}_{1}^{N}\notin\operatorname{ext}\operatorname{\overline{co}}(\hat{X}^{N})$. This already yields ˜5.4.
Alternatively, by the induction hypothesis and upon relabeling tokens if necessary, we have $\{\hat{x}_{1}^{N}\}\cap\hat{X}^{j}=\{\hat{x}_{1}^{N-1}\}\subset\hat{X}^{N-1}$ for all $j\in[N-1]$. Because the sequences are assumed pairwise distinct, one of two possibilities arises.
@@
    There exists $\hat{x}_{2}^{N}\in\hat{X}^{N}$ such that $\hat{x}_{2}^{N}\notin\hat{X}^{N-1}$. In this case, we can apply the same construction of Case 1 or Case 2 to the token $\hat{x}_{2}^{N}$, yielding ˜5.4 with a single-block transformer $\operatorname{\mathrm{T}}^{\prime}$ of width $d^{\prime}\leq 4$.
@@
    There exists $\hat{x}_{2}^{N-1}\in\hat{X}^{N-1}$ such that $\hat{x}_{2}^{N-1}\notin\hat{X}^{N}$. In this case, we first apply Case 1 or Case 2 to $\hat{x}_{2}^{N-1}$ to obtain a single-block transformer $\tilde{\operatorname{\mathrm{T}}}$ such that $\tilde{\operatorname{\mathrm{T}}}_{1}(\hat{X}^{N})\neq\tilde{\operatorname{\mathrm{T}}}_{1}(\hat{X}^{N-1})$. Then, apply again Case 1 or Case 2 to the token $\tilde{\operatorname{\mathrm{T}}}_{1}(\hat{X}^{N})$ to obtain a transformer $\operatorname{\mathrm{T}}^{\prime}$ with two blocks in total and width $d^{\prime}\leq 4$ satisfying ˜5.4.
@@
5.3. Leader Selection
We slightly abuse notation and set $\hat{X}\coloneqq\operatorname{\mathrm{T}}_{\text{sep}}(X)$. The overall goal of this step is to select $m^{j}$ tokens in each sequence to act as leaders, and position them in a suitable way such that, in the next step below, we will be able to collapse the remaining $n^{j}-m^{j}$ tokens and obtain a sequence with $m^{j}$ unique tokens, as required by ˜1.2. We proceed as follows.
@@ 5.3. Leader Selection
the line spanned by $w$. Moreover, choose distinct points $\{x^{1},\dots,x^{N}\}\subset\mathcal{P}\cap Q$ and a hypersphere $S_{R}$ of arbitrary radius $R>0$ (see Figure˜7 for an illustration of all objects).
Figure 7. Sketch of the geometric construction in $\mathbb{R}^{2}$ used in the leader selection step of the proof of Theorem˜1.4. In the left panel, all tokens from the $N=3$ sequences $X^{1}$ (stars), $X^{2}$ (circles) and $X^{3}$ (triangles) lie in the hypercube $\mathcal{B}$. In the right panel, after the leader selection step, $m^{1}=3$, $m^{2}=1$ and $m^{3}=2$ leaders have been selected for each sequence, respectively. For each sequence, one leader is placed in $\mathcal{P}$ and the remaining ones (if any) are placed on $S_{R}\cap\mathbb{R}^{d}_{<0}$.
For every $j\in[N]$, we use Lemma˜4.8 iteratively to obtain:
@@
5.4. Collapse
In this step, we use the clustering properties of the SA layer to collapse all tokens in the box $\mathcal{B}$ to $x^{1},\dots,x^{N}$, while keeping the rest fixed. This is obtained by Lemma˜4.4 with a single-block transformer $\operatorname{\mathrm{T}}_{\text{col}}$ where the FF layer acts as the identity ($\eta=1$, $W=0$, $U=0$, $b=0$) and the parameters in the SA layer are set to $\rho=0$, $V=I_{d}$, and $A=I_{d}$. We thus conclude that $(\operatorname{\mathrm{T}}_{\text{col}}\circ\operatorname{\mathrm{T}}_{\text{lead}}\circ\operatorname{\mathrm{T}}_{\text{sep}})(X^{j})$ has $m^{j}$ unique elements for all $j\in[N]$.
5.5. Interpolation
The transformer $\operatorname{\mathrm{T}}_{\text{col}}\circ\operatorname{\mathrm{T}}_{\text{lead}}\circ\operatorname{\mathrm{T}}_{\text{sep}}$ constructed so far has collapsed the $n^{j}$ distinct tokens of each input sequence $X^{j}$ into $m^{j}$ distinct tokens, as required by the output sequence $Y^{j}$. The remaining task is to move each token iteratively to its desired output. For every $j\in[N]$, we use Lemma˜4.8 to obtain $m^{j}$ FF layers that sequentially map the $m^{j}$ unique tokens of the collapsed sequence $(\operatorname{\mathrm{T}}_{\text{col}}\circ\operatorname{\mathrm{T}}_{\text{lead}}\circ\operatorname{\mathrm{T}}_{\text{sep}})(X^{j})$ to the $m^{j}$ outputs $y_{1}^{j},\dots,y_{m^{j}}^{j}$. This can be implemented by a transformer $\operatorname{\mathrm{T}}_{\text{int}}$ with at most $\sum_{j=1}^{N}m^{j}$ blocks, where each block contains one FF layer provided by Lemma˜4.8 together with a SA layer acting as the identity ($\rho=1$, $V=0$, $A=0$).
5.6. Overall Construction and Complexity Estimate

### v3

@@ 5.1 Proof strategy
Our proof proceeds in four steps, illustrated in Figure 5 for a simple sequence interpolation problem in $\mathbb{R}^{2}$ with $N=3$ input sequences $X^{1}$, $X^{2}$ and $X^{3}$ of length $n^{1}=5$, $n^{2}=3$ and $n^{3}=3$ and output sequences of length $m^{1}=1,m^{2}=1$, and $m^{3}=2$. The initial configuration of the sequences and the desired output sequences are shown in panel $(a)$ of Figure 5.
@@
Figure 5: Proof sketch of Theorem 1.6 applied to $N=3$ input sequences in $\mathbb{R}^{2}$ (denoted with blue circles, orange triangles and green stars, respectively), with output sequences of length $m^{1}=1,m^{2}=1$, and $m^{3}=2$ (denoted with diamonds colored accordingly). The tokens of the initial sequences are shown in panel $(a)$. Panels (b)(b)$(e)$ show the tokens of each sequence after the separation, leader selection, collapse and interpolation steps of the proof. Note that, after the collapse step in panel $(d)$, only tokens selected as leaders are visible. After all steps, sequences match their outputs, as shown in panel $(e)$.
  1. 1.
@@
    Separation. In this step, we use $\mathcal{O}(N)$ transformer blocks to separate overlapping tokens from different input sequences (see panel $(b)$ in Figure 5). This step requires alternating feed-forward and self-attention layers.
  2. 2.
@@
    Leader selection. In this step, we move all tokens to the positive orthant. Then, we select $m^{j}$ points from each sequence to act as “leaders”, and position them carefully using $\mathcal{O}(\sum_{j=1}^{N}m^{j})$ transformer blocks (see panel $(c)$ in Figure 5).
  3. 3.
@@
    Collapse. In this step, all tokens not selected to be leaders collapse onto the leaders by leveraging the clustering property of a single transformer block (see panel $(d)$ in Figure 5).
  4. 4.
@@
    Interpolation. In this final step, the collapsed sequences are driven sequentially to their outputs (see panel $(e)$ in Figure 5), which requires at most $\mathcal{O}(\sum_{j=1}^{N}m^{j})$ transformer blocks.
@@
The next subsections make these steps rigorous to prove Theorem 1.6. Throughout, we use the crucial fact that the ReLU is zero in a half-space, allowing us to move some tokens while leaving the other ones fixed.
5.2 Separation
@@
We begin by ensuring that all input tokens are mapped to non-zero vectors, which can be easily obtained with a single transformer block globally shifting all tokens (setting $U=0$, $b=1$ and $W=w\in\mathbb{R}^{d}$ for any non-zero vector in the FF layer, and $\rho=1$, $V=0$, and $A=0$ in the SA layer). After this initial block, we prove the result constructively by induction on $N$ assuming non-zero tokens, using FF layer constructed as in Lemma 4.7. Throughout, $x_{i}^{j}$ denotes the $i$-th token of sequence $X^{j}$.
@@
$\boldsymbol{N=2}$. By relabeling tokens if necessary, we may assume without loss of generality that $x_{1}^{2}\in X^{2}\setminus X^{1}$, because the sequences are assumed pairwise distinct. We consider two cases, each of which can be handled using one transformer block.
@@
  • Case 1: $x_{1}^{2}\in\operatorname{ext}\operatorname{\overline{co}}(X^{2})$. In this case, the extreme token $x_{1}^{2}\in X^{2}$ allows for sequence separation using a single transformer block of width $d^{\prime}=1$. For the $\operatorname{\mathrm{FF}}$ layer of this block, we apply Lemma 4.7 with $i^{\star}=1$, $j^{\star}=2$ to construct a feed-forward layer $\mathbb{FF}$ and a vector $v\in\mathbb{R}^{d}$ such that, for some $i_{1}\in[m^{1}]$,
    $\mathcal{C}_{r}(\mathbb{FF}(X^{1}),vv^{\top})=\{i_{1}\}\quad\forall r\in[m^{1}],$ (5.1a)
    $\mathcal{C}_{\ell}(\mathbb{FF}(X^{2}),vv^{\top})=\{1\}\quad\forall\ell\in[m^{2}].$ (5.1b)
@@
    Note that $x_{i_{1}}^{1}\in\operatorname{ext}\operatorname{\overline{co}}(X^{1})$ by Lemma 4.5, and $x_{i_{1}}^{1}\neq x_{1}^{2}$ by construction. For the SA layer, instead, we take $0<\alpha<1$ to be specified precisely below and set $\rho=1-\alpha$, $V=\alpha I$ and $A=vv^{\top}$ to obtain
    $\operatorname{\mathrm{SA}}_{r}(\mathbb{FF}(X^{1}))$ $={x}_{r}^{1}+\alpha({x}_{i_{1}}^{1}-{x}_{r}^{1})+\alpha b\frac{v}{\|v\|^{2}}\quad\forall r\in[m^{1}]$
    $\operatorname{\mathrm{SA}}_{\ell}(\mathbb{FF}(X^{2}))$ $={x}_{\ell}^{2}+\alpha({x}_{1}^{2}-{x}_{\ell}^{2})+\alpha b\frac{v}{\|v\|^{2}}\quad\forall\ell\in[m^{2}].$
@@
    (The last term in these two equations comes from the shift introduced by the FF layer.) These quantities are different for all $(r,\ell)\in[m^{1}]\times[m^{2}]$ if and only if
    $(1-\alpha)(x_{\ell}^{2}-x_{r}^{1})+\alpha(x_{1}^{2}-x_{i_{1}}^{1})\neq 0\quad(r,\ell)\in[m^{1}]\times[m^{2}].$ (5.2)
@@
    This condition fails if and only if there exists one pair $(r,\ell)$ and $c_{r\ell}\in\mathbb{R}$ such that $x_{\ell}^{2}-x_{r}^{1}=c_{r\ell}(x_{1}^{2}-x_{i_{1}}^{1})$ and $\alpha$ is the unique solution to $\frac{\alpha}{\alpha-1}=c_{r\ell}$. This means there are at most finitely many choices of $\alpha$ for which (5.2) fails, so it suffices to pick $\alpha$ not from this set.
@@
    Figure 6: Illustration of Case 2 in Lemma 5.1 for $N=2$ sequences in $\mathbb{R}^{2}$. The initial tokens in $X^{1}$ (circles) and $X^{2}$ (stars) are shown in panel $(a)$, as well as the $\operatorname{\mathrm{FF}}$ layer that globally shifts by $v\in\mathbb{R}^{2}$ and moves the distinct token from $X^{2}$ (marked with a red arrow). In panel $(b)$, we show the result of applying the resulting FF layer of width $d^{\prime}=4$. Note how the shifted token is now extreme for $\mathbb{FF}(X^{2})$. In panel $(c)$, the sequences have been separated by the $\operatorname{\mathrm{SA}}$ layer.
@@
    Fix any $y\notin(\operatorname{\overline{co}}(X^{2})\cup X^{1})$. By Lemma 4.8, there exists a width $d^{\prime}=3$ feed-forward layer $\tilde{\operatorname{\mathrm{FF}}}$ such that $\tilde{\operatorname{\mathrm{FF}}}(x_{1}^{2})=y$, and $\tilde{\operatorname{\mathrm{FF}}}(x)=x$ for all other tokens $x\neq x_{1}^{2}$. Using Lemma 4.9 to concatenate $\tilde{\operatorname{\mathrm{FF}}}$ to the width $d^{\prime}=1$ feed-forward layer of Lemma 4.7, we obtain a width $d^{\prime}=4$ feed-forward layer $\mathbb{FF}$ such that 5.1 holds and
    $\operatorname{\mathrm{FF}}(x_{1}^{2})\in\operatorname{ext}\operatorname{\overline{co}}\mathbb{FF}(X^{2}).$ (5.3)
    This inclusion is true because the extreme points of a set are preserved when shifting the set. We can then construct the SA layer in this transformer block following the same construction as in Case 1. The action of the overall transformer block is illustrated in Figure 6.
Induction step. Assume there exists a hardmax transformer $\operatorname{\mathrm{T}}^{N-1}$ with width $d^{\prime}\leq 4$ and at most 1+2(N2)1+2(N-2) blocks such that
@@
Set $\hat{X}^{j}\coloneqq\operatorname{\mathrm{T}}^{N-1}(X^{j})$ for all $j$ to ease the notation. Here and throughout, we write $\operatorname{\mathrm{T}}_{i}(X)$ for the $i$-th token of the transformed sequence $\operatorname{\mathrm{T}}(X)=\{\operatorname{\mathrm{T}}_{1}(X),\dots,\operatorname{\mathrm{T}}_{\operatorname{len}(X)}(X)\}$. Fix without loss of generality $\hat{x}_{1}^{N}\in\hat{X}^{N}$. Our next objective is to build a hardmax transformer $\operatorname{\mathrm{T}}^{\prime}$ with at most two blocks such that
$\operatorname{\mathrm{T}}^{\prime}(\hat{X}^{N})\cap\operatorname{\mathrm{T}}^{\prime}(\hat{X}^{j})=\emptyset\quad\text{for all}\;j\in[N-1].$ (5.4)
If $\{\hat{x}_{1}^{N}\}\cap\hat{X}^{j}=\emptyset$ for all $j\in[N-1]$, then we apply exactly the same argument as in Case 1 or Case 2 above for $N=2$, depending on whether $\hat{x}_{1}^{N}\in\operatorname{ext}\operatorname{\overline{co}}(\hat{X}^{N})$ or $\hat{x}_{1}^{N}\notin\operatorname{ext}\operatorname{\overline{co}}(\hat{X}^{N})$. This already yields 5.4.
If $\{\hat{x}_{1}^{N}\}\cap\hat{X}^{j}\neq\emptyset$, instead, we can use the induction hypothesis to obtain (relabeling tokens and the first N1N-1 sequences if necessary) $\{\hat{x}_{1}^{N}\}\cap\hat{X}^{N-1}=\{\hat{x}_{1}^{N-1}\}\subset\hat{X}^{N-1}$. Since the sequences are assumed pairwise distinct, one of two possibilities arises.
@@
    There exists $\hat{x}_{2}^{N}\in\hat{X}^{N}$ such that $\hat{x}_{2}^{N}\notin\hat{X}^{N-1}$. In this case, we can apply the same construction of Case 1 or Case 2 to the token $\hat{x}_{2}^{N}$, yielding 5.4 with a single-block transformer $\operatorname{\mathrm{T}}^{\prime}$ of width $d^{\prime}\leq 4$.
@@
    There exists $\hat{x}_{2}^{N-1}\in\hat{X}^{N-1}$ such that $\hat{x}_{2}^{N-1}\notin\hat{X}^{N}$. In this case, we first apply Case 1 or Case 2 to $\hat{x}_{2}^{N-1}$ to obtain a single-block transformer $\tilde{\operatorname{\mathrm{T}}}$ such that $\tilde{\operatorname{\mathrm{T}}}_{1}(\hat{X}^{N})\neq\tilde{\operatorname{\mathrm{T}}}_{1}(\hat{X}^{N-1})$. Then, apply again Case 1 or Case 2 to the token $\tilde{\operatorname{\mathrm{T}}}_{1}(\hat{X}^{N})$ to obtain a transformer $\operatorname{\mathrm{T}}^{\prime}$ with two blocks in total and width $d^{\prime}\leq 4$ satisfying 5.4.
@@
5.3 Leader selection
We slightly abuse notation and set $\hat{X}\coloneqq\operatorname{\mathrm{T}}_{\text{sep}}(X)$. The overall goal of this step is to select $m^{j}$ tokens in each sequence to act as leaders, and position them in a suitable way such that, in the next step below, we will be able to collapse the remaining $n^{j}-m^{j}$ tokens and obtain a sequence with $m^{j}$ unique tokens, as required by 1.2. We proceed as follows.
@@ 5.3 Leader selection
the line spanned by $w$. Moreover, choose distinct points $\{x^{1},\dots,x^{N}\}\subset\mathcal{P}\cap Q$ and a hypersphere $S_{R}$ of arbitrary radius $R>0$ (see Figure 7 for an illustration of all objects).
Figure 7: Sketch of the geometric construction in $\mathbb{R}^{2}$ used in the leader selection step of the proof of Theorem 1.6. In the left panel, all tokens from the $N=3$ sequences $X^{1}$ (stars), $X^{2}$ (circles) and $X^{3}$ (triangles) lie in the hypercube $\mathcal{B}$. In the right panel, after the leader selection step, $m^{1}=3$, $m^{2}=1$ and $m^{3}=2$ leaders have been selected for each sequence, respectively. For each sequence, one leader is placed in $\mathcal{P}$ and the remaining ones (if any) are placed on $S_{R}\cap\mathbb{R}^{d}_{<0}$.
For every $j\in[N]$, we use Lemma 4.8 iteratively to obtain:
@@
5.4 Collapse
In this step, we use the clustering properties of the SA layer to collapse all tokens in the box $\mathcal{B}$ to $x^{1},\dots,x^{N}$, while keeping the rest fixed. This is obtained by Lemma 4.4 with a single-block transformer $\operatorname{\mathrm{T}}_{\text{col}}$ where the FF layer acts as the identity ($\eta=1$, $W=0$, $U=0$, $b=0$) and the parameters in the SA layer are set to $\rho=0$, $V=I_{d}$, and $A=I_{d}$. We thus conclude that $(\operatorname{\mathrm{T}}_{\text{col}}\circ\operatorname{\mathrm{T}}_{\text{lead}}\circ\operatorname{\mathrm{T}}_{\text{sep}})(X^{j})$ has $m^{j}$ unique elements for all $j\in[N]$.
5.5 Interpolation
The transformer $\operatorname{\mathrm{T}}_{\text{col}}\circ\operatorname{\mathrm{T}}_{\text{lead}}\circ\operatorname{\mathrm{T}}_{\text{sep}}$ constructed so far has collapsed the $n^{j}$ distinct tokens of each input sequence $X^{j}$ into $m^{j}$ distinct tokens, as required by the output sequence $Y^{j}$. The remaining task is to move each token iteratively to its desired output. For every $j\in[N]$, we use Lemma 4.8 to obtain $m^{j}$ FF layers that sequentially map the $m^{j}$ unique tokens of the collapsed sequence $(\operatorname{\mathrm{T}}_{\text{col}}\circ\operatorname{\mathrm{T}}_{\text{lead}}\circ\operatorname{\mathrm{T}}_{\text{sep}})(X^{j})$ to the $m^{j}$ outputs $y_{1}^{j},\dots,y_{m^{j}}^{j}$. This can be implemented by a transformer $\operatorname{\mathrm{T}}_{\text{int}}$ with at most $\sum_{j=1}^{N}m^{j}$ blocks, where each block contains one FF layer provided by Lemma 4.8 together with a SA layer acting as the identity ($\rho=1$, $V=0$, $A=0$).
5.6 Overall construction and complexity estimate